An inductor, L, of 2 H and a resistor, R, of 1000Ω are in series across an a.c. voltage of 10 V, frequency 60 Hz. Calculate
(i) the impedance in the circuit

(ii) the current in r.m.s. flowing,

(iii) the r.m.s voltage across the inductor.
1
Expert's answer
2020-07-20T14:55:54-0400

(i) The impedance of a series circuit is:


"Z = \sqrt{R^2 + (2\pi fL)^2} = \sqrt{1000^2 + (2\pi \cdot 60\cdot 2)^2} \approx1252.4\Omega"

(ii) the current in r.m.s. flowing, according to the Ohm's law:


"I = \dfrac{V}{Z} = \dfrac{10V}{1252.4\Omega} \approx 0.00798A =7.98mA"

(iii) the r.m.s voltage across the inductor according to the Ohm's law:


"V_L = IZ_L = I\cdot 2\pi fL = 0.00798\cdot 2\pi\cdot 60\cdot 2 \approx 6.02V"

Answer. (i)1252.4 Ohm, (ii)7.98 mA, (iii)6.02 V.